OpenStudy (anonymous):

Help. Question will be showed after i post this. An animal's basal metabolic rate, or BMR, is a measure of the amount of energy the animal uses per day when at rest. An animal's BMR is related to its weight. The more the animal weighs the higher the BMR and the more food it needs to consume in a day. The function f(x)=70x^(3/4) can be used to estimate the BMR of mammals an many other tyes of animals. In this function, x is the mass of the animal in kilograms, and f(x) is its BMR in kilocalories (kcal) per day. Note that 1 kcal is equal to 1 food calorie.

OpenStudy (anonymous):

The howler monkeys at a zoo are fed a diet that contains 3.35 kcal/gram. Write a function g(x) that gives the mass of food in grams that a howler monkey must eat to obtain x kilocalories of energy.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

@satellite73

OpenStudy (aum):

3.35 kcals are obtained from 1 gram of food. x kcals are obtained from x / 3.35 grams of food. \[\large g(x) = \frac{x}{3.35}\]

OpenStudy (aum):

x is kilocalories needed and g(x) is the mass of the food in grams.

OpenStudy (aum):

gtg.

OpenStudy (anonymous):

I think i understand what you mean. So for the question that is continuing off of that one... it asks me to find g(f(x)) so would this be an appropriote response? g(f(x))=70x^(3/4)/3.35

OpenStudy (anonymous):

ok thanks for helping

OpenStudy (aum):

Correct. \[\large f(x) = 70x^{\frac 34} \\ \large g(x) = \frac{x}{3.35} \\ \large g(f(x)) = \frac{70x^{\frac 34}}{3.35} \\ \large \]